Sin(x-y)=cos(x+y)=1/2 
=Sin(x-y)=1/2
=Sin(x-y)=30°
=(X-y)=30°----(1)
Cos(x+y)=1/2
=Cos(x+y)=60°
=(X+y)=60°-----(2)
1or2----
(X-y)=30°
(X+y)=60°

Grade:12th pass

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

Please attach the image of the question

Regards
Abdul Rahman
3 years ago

sin(x-y)=1/2

sin(x-y)= sin300

x-y =300--------1

and cos(x+y) =1/2

       cos(x+y) = cos600

         x+y = 600----------2

on adding 1 & 2, we get,

2x = 900

x = 450

and on subtracting 1 & 2, we get,

-2y = -300

 y = 150